# Accuracy and stability of the accelerated multi-direct-forcing immersed boundary method

### v2

@@ 4.1.3 Sphere fixed in a planar Poiseuille flow
Figure 8: Maximum error of the no-slip condition on a sphere with the weighting function $\phi_{4}$: (a) dependence of the iteration count $\ell$ for various values of the acceleration parameter $\omega$; (b) dependence of $\omega$ without iteration ($\ell=1$).
@@
Fig. 8 shows the maximum error of the no-slip condition on the sphere in steady state. From Fig. 8(a), we can see the same trend as for the two-dimensional results in section 4.1.1 and section 4.1.2, i.e., the error decreases monotonically with the iteration count, and the error for $\omega=C_{4}^{-1}$ is about ten times as small as that for $\omega=1.0$ with $\ell=1$. Also, from Fig. 7(b), we can see that the optimal value of $\omega$ is approximately equal to $C_{4}^{-1}$, in the same way as for the two-dimensional results. Therefore, even for three-dimensional problems, the choice of $\omega=C_{s}^{-1}$ can optimize the accuracy of the no-slip condition. As before, the weighting function $\phi_{3}$ gives similar results (not shown).
@@ 4.2.1 Circular cylinder moving in a Poiseuille flow
Figure 9: Maximum error of the no-slip condition on a moving circular cylinder: (a) dependence of the iteration count $\ell$ for various values of the acceleration parameter $\omega$; (b) dependence of $\omega$ without iteration ($\ell=1$).
First, we check the accuracy of the no-slip condition. We run stable simulations with $D=20\Delta{x}$ and $\gamma=1.0$ until the non-dimensional time $t^{*}=tU/H=50$, by which the linear and angular velocities of the circular cylinder remain almost unchanged. Since the error value of the no-slip condition oscillates with time even in the steady state, we calculate the time-averaged error over $40\leq t^{*}\leq 50$. Fig. 9 shows the maximum error of the no-slip condition on the moving circular cylinder. In the same way as for stationary boundary problems, the error decreases monotonically with the iteration count, and the optimal value of $\omega$ is approximately equal to $C_{4}^{-1}$. Thus, the general behavior of the accuracy of the no-slip condition is independent of stationary or moving boundaries.
@@
It is expected that, in the case with iteration, the numerical stability should be determined by $\eta A$, rather than by $A$, as discussed in section 3.3. Fig. 10(b) shows the results with iteration ($\ell=6$), using $\eta A$ instead of $A$ for the vertical axis. From this figure, we can see that there is an upper limit $\eta A\lessapprox 1.0$ for stable results, and the limit value is almost independent of $\omega$, $\gamma$, and $\Delta{x}/D$ and close to that in the case without iteration.
@@ 4.3.1 Butterfly flight
We consider the same problem as Ref. [23], i.e., free-flight simulations of a butterfly model with wing-pitch flexibility and thorax-pitch control. The system of the butterfly model is a four-body problem composed of two rigid flat plates (left and right wings) and two rigid rods (the thorax and abdomen). We set the maximum thorax-pitch angle to $\alpha_{\rm m}=60^{\circ}$ and the non-dimensional wing-pitch spring stiffness to $N_{\rm K}=2.00$ as one of the computational parameter sets investigated in Ref. [23]. The reference length is given by $L_{\rm ref}=120\Delta{x}$, the flapping period is given by $T=12000\Delta{t}$, and the mean flapping speed is given by $U_{\rm ref}=0.0314\Delta{x}/\Delta{t}$. The boundary points are almost evenly distributed on the model, and the number of the boundary points is $N=14668$ for one wing, which gives $\Delta{V}_{k}=0.978\Delta{x}^{3}$. We use $\phi_{4}$ as the weighting function. The equations of motion of the butterfly model are described by five independent variables (position coordinates $x$ and $y$, thorax-pitch angle $\theta_{\rm th}$, wing-pitch angle $\theta_{\rm wp}$, and abdominal angle $\theta_{\rm ab}$) and calculated by using the second-order Adams–Bashforth method. For more details of the butterfly model and computational conditions, please see Ref. [23].
We set the acceleration parameter $\omega$ and the iteration count $\ell$ as $(\omega,\ell)=(1.0,1)$, $(1.0,6)$, and $(C_{4}^{-1},1)$. The first set corresponds to the conventional direct-forcing IBM, the second set corresponds to the conventional multi-direct-forcing IBM with six iterations, and the third set corresponds to the accelerated multi-direct-forcing IBM with the optimal acceleration parameter without iteration. Hereafter, we refer to the first one as “Conventional 1,” to the second one as “Conventional 6,” and to the third one as “Accelerated.” From the results shown in the previous sections, it is expected that the accuracy of the no-slip condition is comparable between Conventional 6 and Accelerated. Fig. 13 shows the comparison between Conventional 1 and 6 and Accelerated in terms of the vortex structures, trajectories, and joint angles of the butterfly model. From this figure, we can see that the results for Conventional 6 and Accelerated are almost identical, while Conventional 1 shows a visible difference in the trajectory. Thus, Conventional 6 and Accelerated give almost the same results even in a complex problem with non-spherical multiple rigid bodies, while Conventional 1 includes a visible error due to the low accuracy of the no-slip condition. Since the results for Conventional 6 have been extensively validated as mentioned in Ref. [23], Accelerated also gives reasonable results for this problem.
@@
Table 1 shows the accuracy of the no-slip condition and the computational time. In this table, the values of the no-slip error are shown as its time-averaged values in $1\leq t/T\leq 2$, where we cut off the initial transient data. The computational time is measured in $0\leq t/T\leq 2$. These simulations were performed on 24 cores of a Xeon Silver 4214 (2.2 GHz) using Message Passing Interface (MPI). From this table, we can see that the no-slip errors are comparable between Conventional 6 and Accelerated, and the no-slip error for Conventional 1 is much larger than the errors for Conventional 6 and Accelerated, as expected from the results shown in previous sections (e.g., see Fig. 9(a)). The computational time for Accelerated is reduced from that for Conventional 6 due to the reduction of the proportion of the IBM. This is naturally attributed to the fact that the iteration of the IBM is reduce from $\ell=6$ to $1$. Therefore, Accelerated produces results comparable to Conventional 6 while requiring less computational cost.
Table 1: Mean and maximum errors of the no-slip condition, computational time per time step, and proportion of the LBM, IBM, body motion, and other parts to the total computing time in the simulations of the butterfly flight for the conventional and accelerated multi-direct-forcing IBMs.
@@
No-slip error Mean $2.169\times 10^{-2}$ $2.334\times 10^{-3}$ $2.818\times 10^{-3}$
@@
where $M$ is the mass of the whole model, ${\rm W\hskip-1.99168ptR}=M_{\rm w}/M=0.1$ is the wing-mass ratio, and $N_{\rm M}=M/(\rho_{\rm f}L_{\rm ref}^{3})=5.03$ is the non-dimensional mass. In the simulation using Accelerated, the value of $A$ is $8.836\times 10^{-2}\ll 1$, which is consistent with the fact that the simulations are numerically stable. To find the boundary between stable and unstable simulations for this problem, we change the parameter $A$ by changing the non-dimensional mass $N_{\rm M}$. As a result, the motion of the model results in numerical instability for $A\geq 0.635$ ($N_{\rm M}\leq 0.7$), whereas the motion of the model is numerically stable for $A\leq 0.556$ ($N_{\rm M}\geq 0.8$). Thus, there is an upper limit of $A$ for stable results in the same way as in the previous sections, although its limit value is slightly less than that for a moving sphere. This finding is presumably because the model is composed of multiple rigid bodies with complex shape far from a sphere.
@@ 4.3.2 Ice slurry flow
We consider the same problem as Ref. [24], i.e., ice slurry flow within a circular tube. We calculate the motions of many spherical particles in a fluid and the heat transfer between the particles and the fluid. The particles are kept at low temperature $T_{\rm p}=0$, and the tube wall is kept at high temperature $T_{\rm w}=1$. Thus, we calculate not only the flow field but also the temperature field by using the IB-LBM. The descriptions of the LBM and IBM for the temperature field are omitted here and found in Ref. [25]. In this simulation, we set the Reynolds number $Re=1000$, Grashof number $Gr=0$, density ratio $\gamma=1.0$, ice-packing factor ${\rm IPF}=15.0$%, and diameter ratio ${\rm DR}=0.1$ as one of the computational parameter sets investigated in Ref. [24]. In this case, the particle diameter is $D=15\Delta{x}$, the tube diameter is $H=150\Delta{x}$, the number of the boundary points for one particle is $N=714$ (which gives $\Delta{V}_{k}=0.989\Delta{x}^{3}$), the number of particles is $N_{\rm p}=225$, and the cross-sectional-averaged flow speed is $U_{\rm ref}=0.04\Delta{x}/\Delta{t}$. We use $\phi_{4}$ as the weighting function. The equations of motions of particles are solved in the same way as in the case of a single moving sphere shown in section 4.2.3. For more details of the system and computational conditions, please see Ref. [24].
@@
Table 2: Mean and maximum errors of the no-slip and isothermal conditions, computational time per time step, and proportion of the LBMs for flow and temperature fields, IBMs for flow and temperature fields, body motion, and other parts to the total computing time in the simulations of the ice slurry flow for the conventional and accelerated multi-direct-forcing IBMs.
@@
No-slip error Mean $1.232\times 10^{-3}$ $2.024\times 10^{-4}$ $2.819\times 10^{-4}$
@@
Isothermal error Mean $1.178\times 10^{-4}$ $5.113\times 10^{-5}$ $6.310\times 10^{-5}$
@@
Table 2 shows the accuracy of the no-slip and isothermal conditions and the computational time. It is noted that in this simulation, the isothermal condition on particle surfaces is enforced by the multi-direct-forcing IBM for the temperature field, and it is accelerated in the same way as that for the flow field. Thus, this table includes the errors in not only the no-slip condition but also the isothermal condition. The values of the no-slip and isothermal errors are their time-averaged values in $9\leq tU_{\rm ref}/H\leq 10$, where we cut off the initial transient data. The computational time is measured in $0\leq tU_{\rm ref}/H\leq 10$. These simulations were performed on 24 cores of a Xeon Silver 4214 (2.2 GHz) using MPI. From this table, we can see that the no-slip and isothermal errors are comparable between Conventional 6 and Accelerated, while the errors for Conventional 1 are much larger than the errors for Conventional 6 and Accelerated, which can be expected from the results shown in previous sections (e.g., see Fig. 9(a)). The computational time for Accelerated is significantly reduced from that for Conventional 6 due to the reduction of the proportion of the IBMs for flow and temperature fields. Therefore, Accelerated produces results comparable to Conventional 6 while requiring significantly less computational cost. The low accuracy of the no-slip and isothermal conditions in Conventional 1 does not significantly affect the statistical results shown in Fig. 15, since the interactions among hundreds of particles hide the influence of the accuracy.
@@ 5 Conclusions
This study investigated the accuracy and stability of the accelerated multi-direct-forcing immersed boundary method, which corresponds to the Richardson iteration for solving the matrix equation determining the volume force for rigid-body problems. The optimal acceleration parameter $\omega$ for this method should be related to the characteristics of the matrix $\mathcal{A}$ in the matrix equation, which is determined by the weighting function for the interpolation and spreading procedures. With Peskin’s discrete delta function $\phi_{s}$ (whose support is $s\Delta{x}$ where $\Delta x$ is the grid spacing and $s=4$ or $s=3$), the infinity norm $\left\|\mathcal{A}\right\|_{\infty}$ and the constant $C_{s}$ ($C_{4}=0.375$ for $\phi_{4}$ and $C_{3}=0.50$ for $\phi_{3}$) have been previously proposed as candidates for determining the optimal acceleration parameter.
Our numerical results indicate that these values are very close to each other and the corresponding algorithms result in the same accuracy of the no-slip condition. Our analysis of the discretized equations of body motion in moving boundary problems suggests that a single lumped parameter — combining the acceleration parameter, the solid–fluid density ratio, and the spatial resolution of the immersed body — should determine the numerical stability of the body motion.
First, we numerically examined the accuracy of the no-slip condition through stationary boundary problems, including a circular cylinder, an elliptical cylinder, and a sphere fixed in a planar Poiseuille flow. As a result, setting the acceleration parameter as $\omega=C_{s}^{-1}$ is almost the optimal choice minimizing the error of the no-slip condition, and the error for this choice is about ten times smaller than that for $\omega=1$ without iteration. Notably, this scenario holds independently of the distance between the neighboring boundary points, the shape of the boundary, the spatial dimensionality, and whether the weighting function is $\phi_{4}$ or $\phi_{3}$.

### v3

@@ 4.1.3 Sphere fixed in a planar Poiseuille flow
Figure 9: Maximum boundary-velocity error on a sphere with the weighting function $\phi_{4}$: (a) dependence of the iteration count $\ell$ for various values of the acceleration parameter $\omega$; (b) dependence of $\omega$ without iteration ($\ell=1$).
@@
Fig. 9 shows the maximum boundary-velocity error on the sphere in steady state. From Fig. 9(a), we can see the same trend as for the two-dimensional results in section 4.1.1 and section 4.1.2, i.e., the error decreases monotonically with the iteration count, and the error for $\omega=C_{4}^{-1}$ is about ten times as small as that for $\omega=1.0$ with $\ell=1$. Also, from Fig. 8(b), we can see that the optimal value of $\omega$ is approximately equal to $C_{4}^{-1}$, in the same way as for the two-dimensional results. Therefore, even for three-dimensional problems, the choice of $\omega=C_{s}^{-1}$ can optimize the velocity error of the no-slip condition. As before, the weighting function $\phi_{3}$ gives similar results (not shown).
@@ 4.2.1 Circular cylinder moving in a Poiseuille flow
Figure 10: Maximum boundary-velocity error on a moving circular cylinder: (a) dependence of the iteration count $\ell$ for various values of the acceleration parameter $\omega$; (b) dependence of $\omega$ without iteration ($\ell=1$).
First, we check the velocity error of the no-slip condition. We run stable simulations with $D=20\Delta{x}$ and $\gamma=1.0$ until the non-dimensional time $t^{*}=tU/H=50$, by which the linear and angular velocities of the circular cylinder remain almost unchanged. Since the boundary-velocity error oscillates with time even in the steady state, we calculate the time-averaged error over $40\leq t^{*}\leq 50$. Fig. 10 shows the maximum boundary-velocity error on the moving circular cylinder. In the same way as for stationary boundary problems, the error decreases monotonically with the iteration count, and the optimal value of $\omega$ is approximately equal to $C_{4}^{-1}$. Thus, the general behavior of the boundary-velocity error is independent of stationary or moving boundaries.
@@
Goldstein et al. [19] also suggested a stability limit in their feedback IBM, where the upper limit of the time step $\Delta{t}$ is determined by the feedback coefficients. This stability limit might seem similar to the upper limit of $A$ in the present study, since the acceleration parameter $\omega$ corresponds to the feedback coefficient $\beta$, as mentioned in section 2.3. However, they are entirely different from each other. The stability limit in the feedback IBM [19] is for the numerical stability of the flow field around the boundary, whereas the upper limit of $A$ in the present study is for the numerical stability of the body motion in moving boundary problems involving the fluid–body interaction. Although the numerical stability for moving boundary problems has been discussed in previous studies (e.g., [22, 23]), the solid–fluid density ratio $\gamma$ has typically been used as the indicator for determining whether a simulation becomes stable or unstable. However, the present study demonstrates that the parameter $A$ serves as a more appropriate indicator than the density ratio $\gamma$. Therefore, the critical value of $A$ provides a quantitative guideline for the a priori choice of simulation parameters leading to numerically stable simulations of moving boundary problems.
It is expected that, in the case with iteration, the numerical stability should be determined by $\eta A$, rather than by $A$, as discussed in section 3.3. Fig. 11(b) shows the results with iteration ($\ell=6$), using $\eta A$ instead of $A$ for the vertical axis. From this figure, we can see that there is an upper limit $\eta A\lessapprox 1.0$ for stable results, and the limit value is almost independent of $\omega$, $\gamma$, and $\Delta{x}/D$ and close to that in the case without iteration. We can obtain the same conclusion for any value of $\ell$ (not shown). In the following subsections, we primarily discuss the case without iteration, i.e., $\ell=1$ and thus $\eta=1$ (see B), for which the numerical stability is expected to be determined solely by $\eta A=A$.
@@ 4.3.1 Butterfly flight
We consider the same problem as Ref. [24], i.e., free-flight simulations of a butterfly model with wing-pitch flexibility and thorax-pitch control. The system of the butterfly model is a four-body problem composed of two rigid flat plates (left and right wings) and two rigid rods (the thorax and abdomen). We set the maximum thorax-pitch angle to $\alpha_{\rm m}=60^{\circ}$ and the non-dimensional wing-pitch spring stiffness to $N_{\rm K}=2.00$ as one of the computational parameter sets investigated in Ref. [24]. The reference length is given by $L_{\rm ref}=120\Delta{x}$, the flapping period is given by $T=12000\Delta{t}$, and the mean flapping speed is given by $U_{\rm ref}=0.0314\Delta{x}/\Delta{t}$. The gravitational acceleration is given by $G=9.99\cdot 10^{-7}\Delta{x}/(\Delta t)^{2}$. The boundary points are almost evenly distributed on the model, and the number of the boundary points is $N=14668$ for one wing, which gives $\Delta{V}_{k}=0.978\Delta{x}^{3}$. We use $\phi_{4}$ as the weighting function. The equations of motion of the butterfly model are described by five independent variables (position coordinates $x$ and $y$, thorax-pitch angle $\theta_{\rm th}$, wing-pitch angle $\theta_{\rm wp}$, and abdominal angle $\theta_{\rm ab}$) and calculated by using the second-order Adams–Bashforth method. For more details of the butterfly model and computational conditions, please see Ref. [24].
We set the acceleration parameter $\omega$ and the iteration count $\ell$ as $(\omega,\ell)=(1.0,1)$, $(1.0,6)$, and $(C_{4}^{-1},1)$. The first set corresponds to the conventional direct-forcing IBM, the second set corresponds to the conventional multi-direct-forcing IBM with six iterations, and the third set corresponds to the accelerated multi-direct-forcing IBM with the optimal acceleration parameter without iteration. Hereafter, we refer to the first one as “Conventional 1,” to the second one as “Conventional 6,” and to the third one as “Accelerated.” From the results shown in the previous sections, it is expected that the boundary-velocity error is comparable between Conventional 6 and Accelerated. Fig. 14 shows the comparison between Conventional 1 and 6 and Accelerated in terms of the vortex structures, trajectories, and joint angles of the butterfly model. From this figure, we can see that the results for Conventional 6 and Accelerated are almost identical, while Conventional 1 shows a visible difference in the trajectory. Thus, Conventional 6 and Accelerated give almost the same results even in a complex problem with non-spherical multiple rigid bodies, while Conventional 1 includes a visible error due to the large boundary-velocity error. Since the results for Conventional 6 have been extensively validated as mentioned in Ref. [24], Accelerated also gives reasonable results for this problem.
@@
Table 1 shows the velocity error of the no-slip condition and the computational time. In this table, the values of the boundary-velocity error are shown as its time-averaged values in $1\leq t/T\leq 2$, where we cut off the initial transient data. The computational time is measured in $0\leq t/T\leq 2$. These simulations were performed on 24 cores of a Xeon Silver 4214 (2.2 GHz) using Message Passing Interface (MPI). From this table, we can see that the boundary-velocity errors are comparable between Conventional 6 and Accelerated, and the boundary-velocity error for Conventional 1 is much larger than the errors for Conventional 6 and Accelerated, as expected from the results shown in previous sections (e.g., see Fig. 10(a)). The computational time for Accelerated is reduced from that for Conventional 6 due to the reduction of the proportion of the IBM. This is naturally attributed to the fact that the iteration of the IBM is reduce from $\ell=6$ to $1$. Therefore, Accelerated produces results comparable to Conventional 6 while requiring less computational cost.
Table 1: Mean and maximum boundary-velocity errors, computational time per time step, and proportion of the LBM, IBM, body motion, and other parts to the total computing time in the simulations of the butterfly flight for the conventional and accelerated multi-direct-forcing IBMs.
@@
Boundary-velocity error Mean $2.169\times 10^{-2}$ $2.334\times 10^{-3}$ $2.818\times 10^{-3}$
@@
where $M$ is the mass of the whole model, ${\rm W\hskip-1.99168ptR}=M_{\rm w}/M=0.1$ is the wing-mass ratio, and $N_{\rm M}=M/(\rho_{\rm f}L_{\rm ref}^{3})=5.03$ is the non-dimensional mass. In the simulation using Accelerated, the value of $A$ is $8.836\times 10^{-2}\ll 1$, which is consistent with the fact that the simulations are numerically stable. To find the boundary between stable and unstable simulations for this problem, we change the parameter $A$ by changing the non-dimensional mass $N_{\rm M}$. As a result, the motion of the model results in numerical instability for $A\geq 0.635$ ($N_{\rm M}\leq 0.7$), whereas the motion of the model is numerically stable for $A\leq 0.556$ ($N_{\rm M}\geq 0.8$). Thus, there is an upper limit of $A$ for stable results in the same way as in the previous sections, although its limit value is slightly less than that for a moving sphere. This finding is presumably because the model is composed of multiple rigid bodies with complex shape far from a sphere. A similar phenomenon has been observed in vortex-induced-vibration simulations of a circular cylinder [23], where adding a splitter plate to the cylinder increases the critical density ratio for weak-coupling algorithms.
@@ 4.3.2 Ice slurry flow
We consider the same problem as Ref. [25], i.e., ice slurry flow within a circular tube. We calculate the motions of many spherical particles in a fluid and the heat transfer between the particles and the fluid. The particles are kept at low temperature $T_{\rm p}=0$, and the tube wall is kept at high temperature $T_{\rm w}=1$. Thus, we calculate not only the flow field but also the temperature field by using the IB-LBM. The descriptions of the LBM and IBM for the temperature field are omitted here and found in Ref. [26]. In this simulation, we set the Reynolds number $Re=1000$, Grashof number $Gr=0$ (i.e., gravity is absent), density ratio $\gamma=1.0$, ice-packing factor ${\rm IPF}=15.0$%, and diameter ratio ${\rm DR}=0.1$ as one of the computational parameter sets investigated in Ref. [25]. In this case, the particle diameter is $D=15\Delta{x}$, the tube diameter is $H=150\Delta{x}$, the number of the boundary points for one particle is $N=714$ (which gives $\Delta{V}_{k}=0.989\Delta{x}^{3}$), the number of particles is $N_{\rm p}=225$, and the cross-sectional-averaged flow speed is $U_{\rm ref}=0.04\Delta{x}/\Delta{t}$. We use $\phi_{4}$ as the weighting function. The equations of motions of particles are solved in the same way as in the case of a single moving sphere shown in section 4.2.3, although repulsive forces during particle–particle and particle–wall collisions prevent a particle from penetrating into another particle or the tube wall. For more details of the system and computational conditions, please see Ref. [25].
@@
Table 2: Mean and maximum boundary-velocity and temperature errors of the no-slip and isothermal conditions, computational time per time step, and proportion of the LBMs for flow and temperature fields, IBMs for flow and temperature fields, body motion, and other parts to the total computing time in the simulations of the ice slurry flow for the conventional and accelerated multi-direct-forcing IBMs.
@@
Boundary-velocity error Mean $1.232\times 10^{-3}$ $2.024\times 10^{-4}$ $2.819\times 10^{-4}$
@@
Boundary-temperature error Mean $1.178\times 10^{-4}$ $5.113\times 10^{-5}$ $6.310\times 10^{-5}$
@@
Table 2 shows the velocity and temperature errors of the no-slip and isothermal conditions and the computational time. It is noted that in this simulation, the isothermal condition on particle surfaces is enforced by the multi-direct-forcing IBM for the temperature field, and it is accelerated in the same way as that for the flow field. Thus, this table includes the errors in not only the no-slip condition but also the isothermal condition. The values of boundary-velocity and temperature errors are their time-averaged values in $9\leq tU_{\rm ref}/H\leq 10$, where we cut off the initial transient data. The computational time is measured in $0\leq tU_{\rm ref}/H\leq 10$. These simulations were performed on 24 cores of a Xeon Silver 4214 (2.2 GHz) using MPI. From this table, we can see that boundary-velocity and temperature errors are comparable between Conventional 6 and Accelerated, while the errors for Conventional 1 are much larger than the errors for Conventional 6 and Accelerated, which can be expected from the results shown in previous sections (e.g., see Fig. 10(a)). The computational time for Accelerated is significantly reduced from that for Conventional 6 due to the reduction of the proportion of the IBMs for flow and temperature fields. Therefore, Accelerated produces results comparable to Conventional 6 while requiring significantly less computational cost. The large velocity and temperature errors of the no-slip and isothermal conditions in Conventional 1 does not significantly affect the statistical results shown in Fig. 16, since the interactions among hundreds of particles hide the influence of the error.
@@ 5 Conclusions
This study investigated the boundary-velocity error and stability of the accelerated multi-direct-forcing immersed boundary method, which corresponds to the Richardson iteration for solving the matrix equation determining the volume force for rigid-body problems. The optimal acceleration parameter $\omega$ for this method should be related to the characteristics of the matrix $\mathcal{A}$ in the matrix equation, which is determined by the weighting function for the interpolation and spreading procedures. With Peskin’s discrete delta function $\phi_{s}$ (whose support is $s\Delta{x}$ where $\Delta x$ is the grid spacing and $s=4$ or $s=3$), the infinity norm $\left\|\mathcal{A}\right\|_{\infty}$ and the constant $C_{s}$ ($C_{4}=0.375$ for $\phi_{4}$ and $C_{3}=0.50$ for $\phi_{3}$) have been previously proposed as candidates for determining the optimal acceleration parameter.
Our numerical results indicate that these values are very close to each other and the corresponding algorithms result in the same velocity error of the no-slip condition. Our analysis of the discretized equations of body motion in moving boundary problems suggests that a single lumped parameter — combining the acceleration parameter, the solid–fluid density ratio, and the spatial resolution of the immersed body — should determine the numerical stability of the body motion.
First, we numerically examined the velocity error of the no-slip condition through stationary boundary problems, including a circular cylinder, an elliptical cylinder, and a sphere fixed in a planar Poiseuille flow. As a result, setting the acceleration parameter as $\omega=C_{s}^{-1}$ is almost the optimal choice minimizing the boundary-velocity error, and the error for this choice is about ten times smaller than that for $\omega=1$ without iteration. Notably, this scenario holds independently of the distance between the neighboring boundary points, the shape of the boundary, the spatial dimensionality, and whether the weighting function is $\phi_{4}$ or $\phi_{3}$.